OpenStudy (anonymous):

The networking organization you joined is throwing a party. You are in charge of buying the chips, which cost two dollars and fifty cents per bag and salsa, which costs four dollars per jar. The chips & salsa budget you are given totals $60. The inequality 2.5x + 4y 60 represents the possible combinations of chips (x) and salsa (y) you can buy.

OpenStudy (anonymous):

Which of the following does not represent a solution to the inequality? 10 bags of chips and 2 jars of salsa 20 bags of chips and 2 jars of salsa 14 bags of chips and 5 jars of salsa 18 bags of chips and 6 jars of salsa

OpenStudy (anonymous):

@ranga @undeadknight26

undeadknight26 (undeadknight26):

I dunno sorry.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Is the inequality 2.5x + 4y <= 60 ?

OpenStudy (anonymous):

yes, i plugged in all of the choices and all of them are UNDER 60 except for only one, it doesntmake sense ):

OpenStudy (ranga):

The one that is not under is your answer. :)

OpenStudy (anonymous):

oh wow, haha thank you!

OpenStudy (ranga):

"Which of the following does NOT represent a solution to the inequality?"

OpenStudy (anonymous):

yes i see

OpenStudy (anonymous):

can you help me with another one ?

OpenStudy (ranga):

What was your answer?

OpenStudy (anonymous):

D

OpenStudy (ranga):

Yes!

OpenStudy (ranga):

Go ahead and post the question. I can try.

OpenStudy (anonymous):

ok. A rental moving truck can haul a maximum of 7,500 pounds. Dejuan has boxes that weigh 40 pounds and boxes that weigh 75 pounds. How many of each kind of box can he haul in one load? If x represents the number of 40-pound boxes and y represents the number of 75-pound boxes, which graph correctly shows the solution to the problem?

OpenStudy (anonymous):

i know its a dashed line so its either A or D

OpenStudy (ranga):

The rental truck can haul a maximum of 7500 pounds. That means up to and including 7500 pounds. The inequality is 40x + 75y <= 7500. Since it is "less than or equal to", it has to be a solid line. To plot the line we set it "equal to" 7500 and draw the line. 40x + 75y = 7500 when x = 0, 75y = 7500 or y = 100. The y-intercept is 100. when y= 0, 40x = 7500, or x = 187.5. So x-intercept is 187.5 Which graph has a solid line with x-intercept 187.5, y-intercept 100 and is shaded below?

OpenStudy (anonymous):

Its A, I submitted D >.<

OpenStudy (ranga):

I think it is B.

OpenStudy (anonymous):

it was A

OpenStudy (ranga):

If it can haul a maximum of 7500 pounds, it should include 7500 and the inequality is <= 7500 and so it has to be a solid line and not a dashed line. May be someone else can chime in.

OpenStudy (agent0smith):

B is correct, ranga is right, A is wrong.